# Towards Identifiability of Interventional Stochastic Differential Equations

### v3

@@ Appendix A Proofs of Results
Fokker-Planck yields a formula for second-order moments of the SDE stationary distribution (see for example Chapter 5.5 in Särkkä & Solin (2019)), such that,
@@
Now, if we choose $\epsilon$ small enough to guarantee the above bound is strictly less than 1, then the Lyapunov equation given in Equation (70) has a unique solution. Furthermore, the Lyapunov equation has an integral form (Särkkä & Solin, 2019), which states that the unique positive definite matrix $\omega$ that satisfies $L\omega+\omega L^{T}+QQ^{T}=0$ can be written as
@@
By the change of variable theorem induced by the area theorem (Evans, 2018) we have that $f(Bc_{i})$ has a density, and therefore the set of vectors $\{f(Bc_{i})\}_{i=1}^{r}$ is a.s. linearly independent.
@@ B.1 Linear SDE Recovery
For the true simulated linear SDE, $A$ and $B$ are sampled with iid Gaussian entries, and then normalized to have spectral norm equal to $0.9$. The true decay matrix $D$ has each diagonal entry sampled iid from the uniform distribution on the interval $[1,2]$. The model is trained with the loss given in Equation (8), with the Adam optimizer (Kingma, 2014) with an initial learning rate of $0.005$ and $3000$ iterations. Additionally, due to the non-convexity of the objective, on each training instance we run 100 independent initializations and pick the one with the smallest training error. This excludes the oversampled setting with $k=r*\log(n)$, where we get good performance with only 5 independent initializations.
@@ B.3 Synthetic Nonlinear SDE Generalization
We train with 10 interventions, and evaluate on 20 held-out interventions. Training is done with AdamW (Loshchilov, 2017), with initial learning rate $0.003$ and 50000 iterations. The hyperparameter ranges considered are given in Table 3. All experiments were run on CPU.
@@ B.4 Simulated GRN SDE Generalization
We adapt the neural ODE architecture proposed in Lin et al. (2025). The base SDE model under $I_{k}$ parametrizes the change of gene expression via the drift function as below. $A$ and $B$ are the coefficient matrices encoding the modular graph. $\alpha$ is a scaling vector controlling the rate of non-linear activation in the module, where the activation $\sigma$ is the logistic sigmoid. Further, $\beta$ is a bias term that shifts the activation threshold of the modules. The GRN is extracted as $A\,\text{diag}(\alpha\circ\mathbf{1}_{N})B$. The intervention is modeled as a combination of standard basis vectors, $\sum_{j\in I_{k}}e_{j}$, specifying the overexpression. The model also learns a single diffusion coefficient scalar. Lastly, $M$ is a masking matrix blocking the signals from the intervened genes’ regulators. Altogether, the drift is given by,
@@
where $\sigma_{*}$ denotes the MLP as described above. The model loss is the Sinkhorn divergence (Feydy et al., 2019) applied to the true perturbed distribution and the samples drawn from the learned SDE.
@@
We fit the hyperparameters by testing the ODE architecture in Lin et al. (2025) on the same simulated datasets, and then doing zero-shot transfer of the hyperparameters to the SDE models. For Bicycle, we sweep the hyperparameters in Table 4 and use the package defaults for the remaining values. Experiments are run on an nvidia tesla V100 gpu.

### v4

@@ 7 Conclusion
Acknowledgments
This work was made possible by support from the MacMillan Family and the MacMillan Center for the Study of the Non-Coding Cancer Genome at the New York Genome Center. A.Z. is the Sijbrandij Foundation Quantitative Biology Fellow of the Damon Runyon Cancer Research Foundation (DRQ-26-25).
@@ Appendix A Proofs of Results
Fokker-Planck yields a formula for second-order moments of the SDE stationary distribution (see for example Chapter 5.5 in sarkka2019applied), such that,
@@
Now, if we choose $\epsilon$ small enough to guarantee the above bound is strictly less than 1, then the Lyapunov equation given in Equation (70) has a unique solution. Furthermore, the Lyapunov equation has an integral form (sarkka2019applied), which states that the unique positive definite matrix $\omega$ that satisfies $L\omega+\omega L^{T}+QQ^{T}=0$ can be written as
@@
By the change of variable theorem induced by the area theorem (evans2018measure) we have that $f(Bc_{i})$ has a density, and therefore the set of vectors $\{f(Bc_{i})\}_{i=1}^{r}$ is a.s. linearly independent.
@@ B.1 Linear SDE Recovery
For the true simulated linear SDE, $A$ and $B$ are sampled with iid Gaussian entries, and then normalized to have spectral norm equal to $0.9$. The true decay matrix $D$ has each diagonal entry sampled iid from the uniform distribution on the interval $[1,2]$. The model is trained with the loss given in Equation (8), with the Adam optimizer (kingma2014adam) with an initial learning rate of $0.005$ and $3000$ iterations. Additionally, due to the non-convexity of the objective, on each training instance we run 100 independent initializations and pick the one with the smallest training error. This excludes the oversampled setting with $k=r*\log(n)$, where we get good performance with only 5 independent initializations.
@@ B.3 Synthetic Nonlinear SDE Generalization
We train with 10 interventions, and evaluate on 20 held-out interventions. Training is done with AdamW (loshchilov2017decoupled), with initial learning rate $0.003$ and 50000 iterations. The hyperparameter ranges considered are given in Table 3. All experiments were run on CPU.
@@ B.4 Simulated GRN SDE Generalization
We adapt the neural ODE architecture proposed in lin2025interpretable. The base SDE model under $I_{k}$ parametrizes the change of gene expression via the drift function as below. $A$ and $B$ are the coefficient matrices encoding the modular graph. $\alpha$ is a scaling vector controlling the rate of non-linear activation in the module, where the activation $\sigma$ is the logistic sigmoid. Further, $\beta$ is a bias term that shifts the activation threshold of the modules. The GRN is extracted as $A\,\text{diag}(\alpha\circ\mathbf{1}_{N})B$. The intervention is modeled as a combination of standard basis vectors, $\sum_{j\in I_{k}}e_{j}$, specifying the overexpression. The model also learns a single diffusion coefficient scalar. Lastly, $M$ is a masking matrix blocking the signals from the intervened genes’ regulators. Altogether, the drift is given by,
@@
where $\sigma_{*}$ denotes the MLP as described above. The model loss is the Sinkhorn divergence (feydy2019interpolating) applied to the true perturbed distribution and the samples drawn from the learned SDE.
@@
We fit the hyperparameters by testing the ODE architecture in lin2025interpretable on the same simulated datasets, and then doing zero-shot transfer of the hyperparameters to the SDE models. For Bicycle, we sweep the hyperparameters in Table 4 and use the package defaults for the remaining values. Experiments are run on an nvidia tesla V100 gpu.